# SLIM: Skill Learning with Multiple Critics

### 01 Feb 2024 v1

@@ Abstract
Self-supervised skill learning aims to acquire useful behaviors that leverage the underlying dynamics of the environment. Latent variable models, based on mutual information maximization, have been particularly successful in this task but still struggle in the context of robotic manipulation. As it requires impacting a possibly large set of degrees of freedom composing the environment, mutual information maximization fails alone in producing useful manipulation behaviors. To address this limitation, we introduce SLIM, a multi-critic learning approach for skill discovery with a particular focus on robotic manipulation. Our main insight is that utilizing multiple critics in an actor-critic framework to gracefully combine multiple reward functions leads to a significant improvement in latent-variable skill discovery for robotic manipulation while overcoming possible interference occurring among rewards which hinders convergence to useful skills. Furthermore, in the context of tabletop manipulation, we demonstrate the applicability of our novel skill discovery approach to acquire safe and efficient motor primitives in a hierarchical reinforcement learning fashion and leverage them through planning, surpassing the state-of-the-art approaches for skill discovery by a large margin.
@@ I Introduction
Self-supervised methods for skill discovery have been extensively developed in recent years as they enable robots to acquire reusable and transferable knowledge. This flexibility is crucial in dynamic and unstructured environments where robots encounter variations, uncertainties, and unforeseen events. Instead of engineering explicit rules and behaviors for each individual task, robots can learn from data and experiences, making the learning process scalable. This approach can significantly improve the efficiency, scalability, and versatility of robotic systems.
@@
One popular approach to skill learning utilizes the so-called mutual information maximization objective [1] to derive intrinsic rewards [2]. Commonly, this involves training a latent-variable conditioned policy with reinforcement learning which maximizes the mutual information between the latent variable i.e. skill, given as input to the agent, and the agent’s state [2, 3]. While this formulation has been shown to enable an embodied agent to discover behaviors with respect to changing its own state, as in locomotion [4], it struggles in situations where we desire to discover skills that affect degrees of freedom composing the state space outside the agent’s own state. For example, impacting object states, as is the case in robotic manipulation, would require extensive exploration to discover interaction skills.
@@
Figure 1: Skill Learning wIth Multiple critics. Our approach enables the effective combination of multiple objectives for self-supervised skill discovery in robotic manipulation. We learn dedicated critics per intrinsic reward function which is used during policy improvement by taking a weighted combination of their normalized advantages.
@@
A simple way to tackle this challenging case might be to augment intrinsic rewards with additional components that encourage such interactions within the environment. For example, one can introduce a reaching bonus to reward the end-effector of the considered manipulator to get close to the objects composing the considered scene. Furthermore, ensuring the safety of skill discovery is another critical problem recently introduced in  [5]. Adding this also defines an extra reward component that needs to be carefully combined with other reward terms. In this work, we show that a naive implementation of combining these multiple rewards to obtain meaningful and safe interaction skills typically doesn’t work or, in the best case, would require extensive tuning to derive a weighted combination that elicits the desired behavior. To solve this, we introduce a novel multi-critic [6] approach to self-supervised skill discovery that is simple to implement and requires little to no effort to find the right combination of different rewards for safe and effective robotic manipulation skill discovery.
@@
We demonstrate the applicability of our approach to acquiring safe and efficient motor primitives in a hierarchical reinforcement learning (HRL) fashion. Then, we leverage them for rearrangement and object trajectory following tasks through planning, surpassing the state-of-the-art approaches for skill discovery by a large margin.
In summary, our main contributions are
@@
II Approach
II-A Preliminaries
Skill Discovery encompasses unsupervised approaches to reinforcement learning which enable the acquisition of diverse behaviors of a reinforcement learning agent in its environment without specific task rewards. One main approach to this problem relies on mutual information maximization between a latent variable sampled from a fixed distribution which encodes a skill and states visited by a policy conditioned on this skill [2, 4]. This is usually achieved by variational information maximization [1] by maximizing the following bound:
@@ II-A Preliminaries
$\geq\mathbb{E}_{(S,Z)}[\log q_{\eta}(Z|S)].$ (1)
@@
where $A_{i}$ represents advantage functions for each reward component and $\omega_{i}$ represents weights used to combine these signals. In the seminal multi-critic approach, the authors’ primary focus was on learning one critic at a time during updates. They achieved this by utilizing sparse encoding for the weights based on the task being learned. However, they also conducted preliminary experiments to showcase the effectiveness of using equal weights, demonstrating the ability to interpolate between tasks. We build on this last approach in our method and adapted it to self-supervised skill discovery.
II-B Method
We consider a latent-space augmented Markov Decision Process [9] in the domain of robotic manipulation $\mathcal{M}=\bigl{\langle}\mathcal{S},\mathcal{A},\mathcal{P},\mathcal{R},\mathcal{Z}\bigr{\rangle}$ where $\mathcal{S}$ is the state space with state vectors $s\in\mathbb{R}^{42}$ containing robot joint positions, robot joint velocities, object pose, end-effector pose, object linear velocity, object angular velocity, end-effector linear velocity and end-effector angular velocity. We note here that we consider both cartesian positions and orientations in object and end-effector poses. $\mathcal{A}$ is the action space of actions $a\in\mathbb{R}^{7}$ split into two parts: $a_{\text{arm}}\in[-1,1]^{6}$ corresponding to normalised delta pose of the robot’s end-effector in Cartesian space which is converted to joint torques using operational space control (OSC) [10], and ${a_{\text{gripper}}}\in\{0,1\}$ a Boolean action to open or close the gripper. $\mathcal{P}$ is the transition function defining our environment dynamics, $\mathcal{R}$ is the reward function, and $\mathcal{Z}$ is a continuous latent space representing skills. To enable the discovery of meaningful and safe interaction skills, we define $\mathcal{R}$ as a composite reward function consisting of the following reward components:
@@ II-B Method
We propose to use a multi-critic actor learning architecture [6] with three critics for the above reward functions to learn a latent variable skill-conditioned policy $\pi(a|s,z)$ using PPO [11]. Fig. 1 describes our method and, as far as our knowledge goes, this proposition hasn’t been considered in the context of skill discovery. Specifically, we propose to utilize a fully separate multi-network architecture as it was shown to perform better in [6]. One key component in our implementation is that we learn the value functions for each reward function using their respective reward scales but perform a batch normalization of the advantages computed from each critic before combining them with weights for actor learning. This scheme has the advantages of (i) fostering unperturbed critic learning per reward component and (ii) easing the burden of choosing appropriate weights to ensure contributions from each reward component are well-balanced when updating the policy. Practically, we use equal weights to combine the normalized advantages. In addition, we follow the skill composition scheme from  [5] by selecting a sequence of skills to execute in each episode to allow for learning safe skill composition. The full algorithm is detailed in Algorithm 1.
@@
3:     Sample sequence of skills $(z_{1},z_{2})_{n}$ for rollouts
4:     Collect trajectories using $\pi_{\theta}$ and $(z_{1},z_{2})_{n}$
5:     Update $\phi$ with rollout data to maximize (2)
@@
13:     Update policy $\pi_{\theta}$ with PPO, using (3) with $\omega_{i}=1$
@@ III Experiments
Setup We use a tabletop manipulation environment modeled in the IsaacGym simulator [14] illustrated in Fig. 1. The environment includes a Franka Emika Panda robot, a table, and a 5-cm cube. The robot is mounted on the table and is always initialized in a fixed configuration shown in the side view image in Fig. 1. Meanwhile, the object is initialized at a randomly sampled position within an initialization area of dimension 24 x 24 cm, illustrated with the orange square in Fig. 1. The object’s orientation is also initialized randomly using a uniform distribution over axis angle rotations. Compared to tabletop manipulation setups studied in previous works [15, 16, 7, 17] our setup is more challenging as our initial object poses are sampled from a wider distribution and our action space is larger [^1] These works usually consider the Fetch robotics manipulation environments [18] where the gripper orientation is fixed and object initial orientations are also fixed to be aligned with the gripper. As such the actions only control 3-D cartesian displacements while we control 6-D position and orientation displacements.. In our experiments, we leverage the high level of parallelism enabled by IsaacGym by running 5000 instances of our environment in parallel. Furthermore, for all experiments, we use a 6-D continuous latent skill distribution. Intuitively, these skills correspond to representing position and orientation displacements.
@@ III-A Qualitative evaluation
SLIM vs. SLIM ablations: To better understand the effectiveness of the proposed method, we compare SLIM to various ablated versions. These ablations can be grouped in two groups. The first group considers using the same reward functions as SLIM, (4), (5), (6), but combined into a single reward function (by simple summation) and hence a single critic. Note that this provides all the same reward signals used in SLIM except we perform the weighted combination of the rewards before learning a single critic. Our first ablation, henceforth called SLIM_unnormalized_rewards (a.k.a SLIM_ur) consists of summing up all rewards. In Fig. 1(b), the skill policy rollouts with this method shows some grasping and lifting behavior is learned but the trajectories are less diverse than in SLIM which uses multiple critics. Furthermore, to prevent different reward scales to be determinants of performance differences, we define a second ablation version called SLIM_normalized_rewards (a.k.a SLIM_nr). Here, similar to SLIM, we apply normalization to ensure all reward signals are on similar scales before combining them into a single reward function. The trajectories from this version are visualized in Fig. 1(c) showing less diversity than SLIM and mostly failing to learn grasping and picking.
@@
Overall, our ablations show the importance of each component to obtain diverse yet interactive and safe manipulation behaviors. We observe that the three reward components are necessary and complementary to achieve our desired behaviors. In the first group of ablations, we clearly observe the difficulty with combining these three components using normalized or unnormalized sums due to possible interferences between reward signal while learning a skill-conditioned policy. Utilizing the multi-critic architecture with dedicated criticis per reward component helps to alleviate this problem and stabilize learning. Furthermore, we show with the second ablation group that while the multi-critic scheme helps with combining reward components, an omission of any of the three rewards, (4), (5), (6), hampers the overall result.
III-B Quantitative evaluation
@@
Coverage and Safety We evaluate coverage and safety for SLIM, SLIM ablations, and baselines. Coverage is measured over a 50 x 50 x 50 cm centered region discretized into 125 units of 10cm cubes. We evaluate by rolling out 100 trajectories per method repeated for 4 seeds and visualize the mean and standard deviation of the number of cubes covered by the object during the rollouts. Safety is measured as the ratio of safe states according to the indicator function in (6) encountered during the rollouts. Both measures are visualized as violin plots shown in Fig. 3. We observe that SLIM matches the strongest safety baseline SASD while being significantly superior in coverage to baselines.
@@
Our findings, as displayed in Table I, for these six trajectories, suggest that SLIM-based motor primitives can serve within planning algorithms, offering an inherent level of safety. This opens the door to executing complex trajectories over single or multiple objects while ensuring arbitrary safety criteria.
@@
IV Related Work
IV-A Skill Discovery
IV-B Multi Critic Learning
@@ ACKNOWLEDGMENT
We thank Taeyoon Lee, Younghyo Park, and the Control and Dynamics team at NAVER LABS for their SASD codebase upon which this research project was developed. We also thank Jean-Michel Renders, Paul Jansonnie, and all team members of Robot Learning at NAVER LABS Europe for insightful discussions and suggestions.

### 21 Mar 2024 v2

@@ Abstract
Self-supervised skill learning aims to acquire useful behaviors that leverage the underlying dynamics of the environment. Latent variable models, based on mutual information maximization, have been successful in this task but still struggle in the context of robotic manipulation. As it requires impacting a possibly large set of degrees of freedom composing the environment, mutual information maximization fails alone in producing useful and safe manipulation behaviors. Furthermore, tackling this by augmenting skill discovery rewards with additional rewards through a naive combination might fail to produce desired behaviors. To address this limitation, we introduce SLIM, a multi-critic learning approach for skill discovery with a particular focus on robotic manipulation. Our main insight is that utilizing multiple critics in an actor-critic framework to gracefully combine multiple reward functions leads to a significant improvement in latent-variable skill discovery for robotic manipulation while overcoming possible interference occurring among rewards which hinders convergence to useful skills. Furthermore, in the context of tabletop manipulation, we demonstrate the applicability of our novel skill discovery approach to acquire safe and efficient motor primitives in a hierarchical reinforcement learning fashion and leverage them through planning, significantly surpassing baseline approaches for skill discovery.
@@ I Introduction
Self-supervised methods for skill discovery have been extensively developed in recent years as they enable robots to acquire reusable and transferable knowledge. This flexibility is crucial in dynamic and unstructured environments where robots encounter variations, uncertainties, and unforeseen events. Instead of engineering explicit rules and behaviors for each individual task, robots can learn from data and experiences, making the learning process scalable, thus improving the efficiency and versatility of robotic systems.
@@
One popular approach to skill discovery utilizes the so-called mutual information maximization objective [1] to derive intrinsic rewards [2]. Commonly, this involves training a latent-variable conditioned policy with reinforcement learning which maximizes the mutual information between the latent variable i.e. skill, given as input to the agent, and the agent’s state [2, 3]. While this formulation has been shown to enable an embodied agent to discover behaviors with respect to changing its own state, as in locomotion [4], it struggles in situations where we desire to discover skills that affect degrees of freedom composing the state space outside the agent’s own state. For example, impacting object states, as is the case in robotic manipulation, would require extensive exploration to discover interaction skills.
@@
Figure 1: Skill Learning wIth Multiple critics. Our approach enables the effective combination of multiple objectives for self-supervised skill discovery in robotic manipulation. We learn dedicated critics per intrinsic reward function which is used during policy improvement by taking a weighted combination of their normalized advantages. (a) Schematic diagram (b) Simulation top view (c) Simulation side view.
@@
A simple way to tackle this challenging case might be to augment intrinsic rewards with additional components that encourage such interactions within the environment. For example, one can introduce a reaching bonus to reward the end-effector of the considered manipulator to get close to the objects composing the considered scene. Furthermore, ensuring the safety of skill discovery is another critical problem recently introduced in  [5]. Adding this also defines an extra reward component that needs to be carefully combined with other reward terms. In this work, we show that a naive implementation of combining these multiple rewards to obtain meaningful and safe interaction skills typically doesn’t work or, in the best case, would require laborious tuning to derive a weighted combination that elicits the desired behavior. To solve this, we introduce a novel multi-critic [6] approach to self-supervised skill discovery that is simple to implement and requires little to no effort to find the right combination of different rewards for safe and effective robotic manipulation skill discovery.
@@
We demonstrate the applicability of our approach for acquiring safe and effective motor primitives in a hierarchical reinforcement learning (HRL) fashion. Then, we leverage them for rearrangement and object trajectory following tasks through planning, surpassing the state-of-the-art baseline approaches for skill discovery.
In summary, our main contributions are:
@@
II Related Work
II-A Skill Discovery
II-B Multi Critic Learning
III Approach
III-A Preliminaries
Skill Discovery encompasses unsupervised approaches to reinforcement learning which enable the acquisition of diverse behaviors of a reinforcement learning agent in its environment without specific task rewards. One main approach to this problem relies on mutual information maximization between a latent variable sampled from a fixed distribution $p(z)$ which encodes skills and states visited by a policy conditioned on this skill [2, 4]. This is usually achieved by variational information maximization [1], by maximizing the following bound:
@@ III-A Preliminaries
$\geq\mathbb{E}_{(s,z)}[\log q_{\eta}(z|s)],$ (1)
where $z\sim p(z)$ represents skills, $s$ represents states from an agent’s trajectory $\tau=(s_{0},...,s_{T})$, and $q_{\eta}(z|s)$ is a discriminator network approximating the posterior distribution of skills given states.
@@
where $A_{i}$ represents advantage functions for each reward component and $\omega_{i}$ represents weights used to combine these signals. In the seminal multi-critic approach, the authors’ primary focus was on learning one critic at a time during updates. They achieved this by utilizing sparse encoding for the weights based on the task being learned. However, they also conducted preliminary experiments to showcase the effectiveness of using equal weights, demonstrating the ability to interpolate between tasks. We build on this approach in our method and adapt it to self-supervised skill discovery.
III-B Method
We consider a Markov Decision Process [28] augmented with a skill latent space in the domain of robotic manipulation $\mathcal{M}=\bigl{\langle}\mathcal{S},\mathcal{A},\mathcal{P},\mathcal{R},\mathcal{Z}\bigr{\rangle}$, where $\mathcal{S}$ is the state space with state vectors $s\in\mathbb{R}^{42}$ containing robot joint positions, robot joint velocities, object pose, end-effector pose, object linear velocity, object angular velocity, end-effector linear velocity and end-effector angular velocity. We note here that we consider both cartesian positions and orientations in object and end-effector poses. $\mathcal{A}$ is the action space of actions $a\in\mathbb{R}^{7}$ split into two parts: $a_{\text{arm}}\in[-1,1]^{6}$ corresponding to normalised delta pose of the robot’s end-effector in Cartesian space which is converted to joint torques using operational space control (OSC) [29], and ${a_{\text{gripper}}}\in\{0,1\}$ a Boolean action to open or close the gripper. $\mathcal{P}$ is the transition function defining our environment dynamics, $\mathcal{R}$ is the reward function, and $\mathcal{Z}$ is a continuous latent space representing skills. To enable the discovery of meaningful and safe interaction skills, we define $\mathcal{R}$ as a composite reward function consisting of the following reward components:
@@ III-B Method
We propose to use a multi-critic actor learning architecture [6] with three critics for the above reward functions to learn a latent variable skill-conditioned policy $\pi(a|s,z)$ using PPO [30]. Fig. 0(a) illustrates our method and, as far as our knowledge goes, this proposition hasn’t been considered in the context of skill discovery. Specifically, we propose to utilize a fully separate multi-network architecture as it was shown to perform better in [6]. One key component in our implementation is that we learn the value functions for each reward function using their respective reward scales but perform a batch normalization of the advantages computed from each critic before combining them with weights for actor learning. This scheme has the advantages of (i) fostering unperturbed critic learning per reward component and (ii) easing the burden of choosing appropriate weights to ensure contributions from each reward component are well-balanced when updating the policy. Practically, we use equal weights to combine the normalized advantages. In addition, we follow the skill composition scheme from  [5] by selecting a sequence of skills to execute in each episode which encourages learning safe skill composition. The full algorithm is detailed in Algorithm 1.
@@
3:     Sample sequence of skills $(z_{1},...,z_{n})$ for rollouts
4:     Collect trajectories using $\pi_{\theta}$ and $(z_{1},...,z_{n})$
5:     Update $\phi$ with rollout data to maximize Eq. (2)
@@
13:     Update $\pi_{\theta}$ with PPO, using Eq. (3) with $\omega_{i}=1$
@@ IV Experiments
Setup We use a tabletop manipulation environment modeled in the IsaacGym simulator [33] illustrated in Fig. 0(b) and Fig. 0(c). The environment includes a Franka Emika Panda robot, a table, and a 5-cm cube. The robot is mounted on the table and is always initialized in a fixed configuration shown in the side view image in Fig. 0(c). Meanwhile, the object is initialized at a randomly sampled position within an initialization area of dimension 24 x 24 cm, illustrated with the orange square in Fig. 0(b). The object’s orientation is also initialized randomly using a uniform distribution over axis angle rotations. Compared to tabletop manipulation setups studied in previous works [16, 17, 14, 15], our setup is more challenging as our initial object poses are sampled from a wider distribution and our action space is larger [^1] These works usually consider the Fetch robotics manipulation environments [34] where the gripper orientation is fixed and object initial orientations are also fixed to be aligned with the gripper. As such the actions only control 3-D cartesian displacements while we control 6-D position and orientation displacements.. In our experiments, we leverage the high level of parallelism enabled by IsaacGym by running 5000 instances of our environment in parallel. Furthermore, for all experiments, we use a 6-D von Mises-Fisher distribution as the fixed prior skill distribution. Intuitively, these skills correspond to representing position and orientation displacements.
@@ IV-A Qualitative evaluation
SLIM vs. SLIM ablations: To better understand the effectiveness of the proposed method, we compare SLIM to various ablated versions. These ablations can be grouped in two groups. The first group considers using the same reward functions as SLIM, Eq. (4), Eq. (5), Eq. (6), but combined into a single reward function (by simple summation) and hence a single critic. Note that this provides all the same reward signals used in SLIM except we perform the weighted combination of the rewards before learning a single critic. Our first ablation, henceforth called SLIM_unnormalized_rewards (a.k.a SLIM_ur) consists of summing up all rewards. In Fig. 1(b), the skill policy rollouts with this method shows some grasping and lifting behavior is learned but the trajectories are less diverse than in SLIM which uses multiple critics. Furthermore, to prevent different reward scales to be determinants of performance differences, we define a second ablation version called SLIM_normalized_rewards (a.k.a SLIM_nr). Here, similar to SLIM, we apply normalization to ensure all reward signals are on similar scales before combining them into a single reward function. The trajectories from this version are visualized in Fig. 1(c) showing less diversity than SLIM and mostly failing to learn grasping and picking.
@@
Overall, our ablations show the importance of each component to obtain diverse yet interactive and safe manipulation behaviors. We observe that the three reward components are necessary and complementary to achieve our desired behaviors. In the first group of ablations, we clearly observe the difficulty with combining these three components using normalized or unnormalized sums due to possible interferences between reward signal while learning a skill-conditioned policy. Utilizing the multi-critic architecture with dedicated criticis per reward component helps to alleviate this problem and stabilize learning. Furthermore, we show with the second ablation group that while the multi-critic scheme helps with combining reward components, an omission of any of the three rewards, Eq. (4), Eq. (5), Eq. (6), hampers the overall result.
IV-B Quantitative evaluation
@@
Coverage and Safety We evaluate coverage and safety for SLIM, SLIM ablations, and baselines. Coverage is measured over a 50 x 50 x 50 cm centered region discretized into 125 units of 10cm cubes. We evaluate by rolling out 100 trajectories per method repeated for 4 seeds and visualize the mean and standard deviation of the number of cubes covered by the object during the rollouts. Safety is measured as the ratio of safe states according to the indicator function in Eq. (6) encountered during the rollouts. Both measures are visualized as violin plots shown in Fig. 3. We observe that SLIM matches the strongest safety baseline SASD while being significantly superior in coverage to baselines.
@@
Our findings, as displayed in Table I, suggest that SLIM-based motor primitives can serve within planning algorithms, offering an inherent level of safety. This opens the door to executing complex trajectories over single or multiple objects while ensuring arbitrary safety criteria.
@@ ACKNOWLEDGMENT
We thank Taeyoon Lee, Younghyo Park, and the Control and Dynamics team at NAVER LABS for their SASD codebase upon which this research project was developed. We also thank Paul Jansonnie and all team members of Robot Learning at NAVER LABS Europe for insightful discussions and suggestions.